# Naturalistic Computational Cognitive Science Towards generalizable models and theories that capture the full range of natural behavior

### v1

@@ 6.2 Incorporating naturalistic variation to enhance generalizability
By testing hypotheses in settings where other aspects of the task and data generating process are sampled broadly from the naturalistic distribution, we address some of the challenges of generalizability that arise from testing on a much narrower distribution than the written hypothesis implies (cf. Yarkoni,, 2022). For example, if the hypothesis is about visual processing in general, but the task paradigm relies solely on abstract shape stimuli, the conclusions may not generalize in the way the written results would suggest. Thus, testing our hypotheses across varied settings—“design[ing] with variability in mind” (Yarkoni,, 2022)—increases the likelihood of identifying robust effects that will generalize (cf. Baribault et al.,, 2018). In the language of regression, varying as many other data generating factors as possible brings us closer to estimating the “main effect” of an experimentally-manipulated construct, rather than estimating the “simple effect” of the manipulation in the single setting we have tested.
For example, it is a long-standing principle of experimental design to test a hypothesis with multiple stimuli, and statistically quantify the stimulus effect to estimate the generalizability (e.g. Clark,, 1973). However, many other experiment features—even exact task formulations such as multi-arm bandit tasks—are often preserved within a paper, or even across an entire literature. Just like testing a hypothesis with a single stimulus, this lack of variation poses challenges for generalizability (cf. Eckstein et al.,, 2022)—especially because the way that researchers operationalize a theory can substantially affect their conclusions (e.g. Holzmeister et al.,, 2024; Schweinsberg et al.,, 2021; Strickland and Suben,, 2012). While naturalistic variation cannot resolve this issue in full, we believe that incorporating a broader range of the natural variation of settings, stimuli, and tasks within which we expect those theories to hold, will help us to develop more generalizable theories.
Incorporating variation into our experiments can also help us to revise the constructs underlying our theories. For example, Eisenberg et al., (2019) study the putative construct of self-regulation by simultaneously testing a large variety of existing self-regulation measures. The authors find that these measures do not form a unitary construct. Rather, task- and survey-based measures each tap into distinct, multi-factorial constructs, with individual tasks loading primarily onto a subset of the factors in their domain. That is, two different studies that each use only a single measure for self-regulation may be tapping into entirely different constructs. By contrast, by substantially varying the tasks we use to measure a theoretical component like self-regulation, we can more fully disentangle the underlying constructs—and thereby improve the extent to which our theories generalize.
@@ 6.3 Interpreting complex models to yield explanations
If we perform experiments in naturalistic settings, and develop complex models that can predict human behavior across them, how does this support explaining the cognitive phenomena of interest? In this section we will review some paths to deriving explanations of phenomena from complex models. [^4] N.B., not all complex systems admit simple, abstract explanations. For example, where natural intelligence incorporates chaotic (Freeman,, 1995) or critical (O’Byrne and Jerbi,, 2022) dynamics, they may alter the kind of explanations we can seek (cf. Kellert,, 1993). A benefit of seeking predictive models as a route to explanation is that the models may be useful even if simple explanations do not follow.
@@
Data properties as explanation & rational analysis. One route to explaining model behavior is as a consequence of the properties of the data from which it learns. This idea stretches back through behaviorism as well as cognitive science; for example, connectionist research has often focused on how naturalistic data could drive cognitive phenomena (McClelland and Jenkins,, 1991; Elman,, 1996; Rogers and McClelland,, 2004). Analogously, in AI, various researchers have taken inspiration from behaviors that emerge in large pretrained models, and similarly attempted to identify minimal properties of naturalistic data that give rise to those behaviors. For example, Chan et al., (2022) explore how the bursty, long-tailed nature of natural data can give rise to in-context learning, and Prystawski et al., (2024) examine how local dependencies can yield certain kinds of sequential reasoning. These works illustrate how rich properties of natural data can be distilled down to the minimal elements that give rise to a behavior—and can thus likewise offer a candidate explanation for the origin of those capabilities. Moreover, explanations of behavior in terms of the property of the data from which a system learns offer a route toward normative explanations of that behavior as a rational solution to constraints—precisely the perspective taken by rational analysis (Anderson,, 1991).
@@
Mechanistic explanation. Understanding that a behavior arises from properties of data does not explain how that behavior is implemented. Fortunately, the internal workings of task-performing computational models can generally be inspected and intervened upon. For example, in contemporary AI research, a variety of works have analyzed the mechanisms that implement model behaviors, either in simplified settings (e.g. Nanda et al.,, 2023; Zhong et al.,, 2024), or even in large models trained on natural data (Geiger et al.,, 2021; Wu et al.,, 2024). A particularly promising approach for cognitive science is illustrated by Distributed Alignment Search (Geiger et al.,, 2024), which attempts to map postulated abstract causal models onto low-level features of the model that play corresponding causal roles—thus enabling the link between abstract models or hypotheses about the computational solution, and evidence about concrete implementational mechanisms. Morever, mechanistic interventions can even be combined with data studies, to examine the causal role of different mechanisms over the course of learning (Singh et al.,, 2024). Appropriate regularization of models can also produce representations that yield greater interpretability—e.g. Miller et al., (2024) train networks to predict animal behavior while restricting their representations, and used these to identify non-optimal features influencing the decisions—thus allowing bottom-up discovery of possible implementations. Thus, various methods for mechanistic study of models can offer powerful routes to experimentally determining sufficient implementations of cognitive processes.
@@
Formal theories. From the way that behavior emerges from data or mechanisms, we can progress to formal theories. Saxe et al., (2019) offers an illustrative example. Prior works observed empirically that pseudo-naturalistic data give rise to various aspects of human-like semantic development and representation in neural networks (Rogers and McClelland,, 2004), such as progressive differentiation. However, these works had not explained why these parallels emerge. Saxe et al., (2019) bridged this gap by formally deriving how the data properties combine with gradient-based learning to yield these cognitive phenomena. This illustrates how preliminary explanations in terms of data can be developed into more rigorous analytic theories. Thus, building naturalistic models can be a stepping stone toward more rigorous, normative theories of intelligence as a rational solution to a particular problem setting—much like prior work in cognitive science (e.g. Anderson,, 1991; Oaksford and Chater,, 2009).
@@
These two components serve different purposes. The first helps to demonstrate that the models are real candidate models of the phenomena in question, rather than models that simply could not scale to the real problem or that are overfit to a particular instantiation of the task. We believe this component is necessary because otherwise the problem space is under-constrained, as we have argued above. Moreover, it provides a test-bed for performing experimentation that would be impossible in reality (e.g. beyond what would be feasible or ethical in the lab), which is useful for both scientific and practical reasons. The second component—where feasible—allows linking these predictions to the explanatory understanding that cognitive scientists have usually sought, and that may help us to generalize our insights beyond the current paradigms. This perspective aligns with past arguments on how deep learning can contribute to understanding in cognitive neuroscience (Saxe et al.,, 2021; Kanwisher et al.,, 2023; Cao and Yamins, 2024b, ; Doerig et al.,, 2023). However, we emphasize that combining these two components makes explicit links between the actual problem the system solves and the theorized constructs underlying that solution. It also ensures that our theories make testable predictions beyond the narrow settings of a particular task paradigm.
@@ 7 Discussion
In this paper, we have tried to outline a direction of research that we call “naturalistic computational cognitive science” that aims to build generalizable models of cognition that scale to naturalistic tasks, while still offering routes to explanatory and theoretical understanding. This perspective synthesizes a growing literature on the importance of naturalistic experiments and generalizable models, and grows out of a long-standing focus in cognitive science on explaining a broad range of phenomena. We outline these connections here.
@@
The quest for building general models of natural intelligence. A long history of cognitive science research has sought to develop frameworks with the generality to explain a wide scope of cognitive phenomena. Researchers in connectionist (e.g. McClelland et al.,, 1986, 2003), Bayesian (e.g. Tenenbaum and Griffiths,, 2001), and cognitive architectures (e.g. Ritter et al.,, 2019; Laird,, 2019) paradigms have tried to identify underlying principles or mechanisms that explain many phenomena. However, while the modeling frameworks themselves are general, typical instantiations of these frameworks have built specialized models focused on individual tasks.
@@
Other recent studies have built upon foundation models from AI that can perform many tasks. Some works have applied cognitive paradigms to study the behaviors of these models (e.g. Binz and Schulz,, 2023; Lampinen et al.,, 2024; Buschoff et al.,, 2025). Other recent works have finetuned these models using cognitive data, to create generalizable cognitive models that can predict behavior on new experimental paradigms in areas like vision (e.g. Muttenthaler et al., 2024b, ; Muttenthaler et al., 2024a, ; Fu et al.,, 2024), or the broader space of cognitive tasks that can be presented in language (Binz and Schulz,, 2024; Binz et al.,, 2024).
@@
We see the arguments that we have laid out in this paper as being broadly compatible with many of these approaches and frameworks, but offering a stronger emphasis on the virtuous cycle of expanding the scope and naturalism of our experimental designs, and expanding the generalizability of our models and theories—and anchoring these in the practicalities of achieving frictionless reproducibility.
@@
Towards experiments that involve naturalistic behavior In emphasizing more naturalistic experiments, our perspective aligns with a variety of works that have likewise highlighted the importance of considering the scope of complex naturalistic behaviors. This perspective has been advocated frequently in neuroscience (e.g. Mobbs et al.,, 2021; Cisek and Green,, 2024), where even prominent visual processing signals can be fundamentally different in active, naturalistic paradigms than in simpler classic ones (e.g. Amme et al.,, 2024). Likewise, Wise et al., (2023) advocate for the need to explore more naturalistic environments and behaviors in reinforcement learning, to grapple with the complexity of behavior in environments closer to the real world. We concur with these works, and emphasize that advances in technology both enable richer experimental paradigms, and the models that can perform them. We differ from many of these works in emphasizing the role of models that generalize across task paradigms.
@@
Linking between learning models and the brain A large recent literature has explored the surprising alignment between the representations learned by task-optimized deep learning models and those in the brain (Khaligh-Razavi and Kriegeskorte,, 2014; Yamins et al.,, 2014; Yamins and DiCarlo,, 2016; Schrimpf et al.,, 2021; Sucholutsky et al.,, 2023). The observations of surprising alignment have led to substantial interest and debate. The reactions range from critical argument that failures to capture particular phenomena doom these models (e.g. Bowers et al.,, 2023), to suggestions that deep learning models upend the theoretical assumptions of broad areas of the field (Perconti and Plebe,, 2020; Hasson et al.,, 2020). Other authors have written frameworks that attempt to integrate deep learning models within the existing scientific paradigms (Doerig et al.,, 2023). In this vein, several papers have proposed frameworks for understanding the role of deep learning in cognitive neuroscience (e.g Richards et al.,, 2019; Storrs and Kriegeskorte,, 2019; Cichy and Kaiser,, 2019)—and how these models can contribute to deriving explanatory and theoretical understanding (Cao and Yamins, 2024a, ; Cao and Yamins, 2024b, ; Saxe et al.,, 2021). In this context, Feather et al., (2025) argue for the importance of benchmarks that place stronger constraints on models, by testing alignment of both behavior and representations. As above, our perspective aligns with many of these works, but we place greater emphasis on the complementary role of increasing naturalism of experimental paradigms, and learning from the engineering paradigms of AI, rather than simply adopting its models as artifacts.

### v2

@@ 6.2 Incorporating naturalistic variation to enhance generalizability
By testing hypotheses in settings where other aspects of the task and data generating process are sampled broadly from the naturalistic distribution, we address some of the challenges of generalizability that arise from testing on a much narrower distribution than the written hypothesis implies (cf. Yarkoni,, 2022). In particular, testing our hypotheses across varied settings—“design[ing] with variability in mind” (Yarkoni,, 2022)—increases the likelihood of identifying robust effects that will generalize (cf. Baribault et al.,, 2018). In the language of regression, varying as many other data generating factors as possible brings us closer to estimating the “main effect” of an experimentally-manipulated construct, rather than estimating the “simple effect” of the manipulation in the single setting we have tested.
For example, it is a long-standing principle of experimental design to test a hypothesis with multiple stimuli, and statistically quantify the stimulus effect to estimate generalizability (e.g. Clark,, 1973). However, many other experiment features—even exact task formulations such as multi-arm bandit tasks—are often preserved within a paper, or even across an entire literature. Just like testing a hypothesis with a single stimulus, this limited variation poses challenges for generalizability (Eckstein et al.,, 2022)—especially because the way that researchers operationalize a theory can substantially affect their conclusions (e.g. Holzmeister et al.,, 2024; Schweinsberg et al.,, 2021; Strickland and Suben,, 2012). While naturalistic variation cannot resolve this issue in full, we believe that incorporating a broader range of the settings, stimuli, and tasks within which we expect a theory to hold, will help us to develop more generalizable theories.
Incorporating variation can also help us to revise the constructs underlying our theories. For example, Eisenberg et al., (2019) study the putative construct of self-regulation by simultaneously testing many existing self-regulation measures. The authors find that these measures do not form a unitary construct. Rather, task- and survey-based measures each form distinct, multi-factorial constructs, with individual tasks loading primarily onto a subset of factors. That is, two different studies that each use only a single measure for self-regulation may be measuring entirely different constructs. By contrast, by substantially varying the tasks we use to measure a theoretical component like self-regulation, we can more fully disentangle the underlying constructs—and thereby improve the extent to which our theories generalize.
@@ 6.3 Interpreting complex models to yield explanations
If we perform experiments in naturalistic settings, and develop complex models that can predict human behavior across them, how does this support explaining the cognitive phenomena of interest? Here, we illustrate some paths to deriving explanations from complex models. [^4] N.B., not all complex systems admit abstract explanations. Where natural intelligence incorporates chaotic (Freeman,, 1995) or critical (O’Byrne and Jerbi,, 2022) dynamics, they may alter the explanations that we can seek (cf. Kellert,, 1993). A benefit of seeking predictive models as a route to explanation is that the models may be useful even if simple explanations do not follow.
@@
Data properties as explanation & rational analysis. One way to explain model behavior is as a consequence of the properties of the data from which it learns. This idea has a rich history in cognitive science andd AI (McClelland and Jenkins,, 1991; Elman,, 1996; Rogers and McClelland,, 2004), in which researchers attempted to identify minimal properties of naturalistic data that give rise to some behavior(s). For example, Chan et al., (2022) explore how the bursty, long-tailed nature of natural data can give rise to in-context learning, and Prystawski et al., (2024) examine how local dependencies can yield certain kinds of sequential reasoning. These works illustrate how rich properties of natural data can be distilled down to minimal elements that drive a behavior—and can thus offer a candidate explanation for the origin of those capabilities. Moreover, explanations of behavior via properties of the data offer a route toward normative explanations of that behavior as a rational solution to constraints—precisely the perspective taken by rational analysis (Anderson,, 1991).
@@
Mechanistic explanation. Understanding that a behavior arises from data properties does not explain how that behavior is implemented. Fortunately, the internal workings of task-performing computational models can generally be inspected and intervened upon. For example, a variety of works have analyzed the mechanisms that implement language model behaviors, either in simplified settings (e.g. Nanda et al.,, 2023; Zhong et al.,, 2024), or in large models trained on natural data (Geiger et al.,, 2021; Wu et al.,, 2024). A particularly promising approach for cognitive science is illustrated by Distributed Alignment Search (Geiger et al.,, 2024), which attempts to map postulated abstract causal models onto low-level features of a complex model that play corresponding causal roles—thus enabling linking between abstract models or hypotheses about the computational solution, and evidence about concrete implementational mechanisms. Morever, mechanistic interventions can even be combined with data studies, to examine the causal role of different mechanisms over the course of learning (Singh et al.,, 2024). Appropriate regularization of models can also produce representations that yield greater interpretability—e.g. Miller et al., (2024) train networks to predict animal behavior while restricting their representations, and used these to identify non-optimal features influencing the decisions—thus allowing bottom-up discovery of possible implementations. Thus, various methods for mechanistic study of models can offer powerful routes to experimentally determining sufficient implementations of cognitive processes.
@@
Formal theories. From data properties or mechanisms, we can progress to formal theories. Saxe et al., (2019) offers an illustrative example. Prior works observed empirically that pseudo-naturalistic data elicit various aspects of human-like semantic development and representation in neural networks (Rogers and McClelland,, 2004), such as progressive differentiation. However, these works had not explained why these parallels emerge. Saxe et al., (2019) bridged this gap by formally deriving how the data properties combine with gradient-based learning to yield these cognitive phenomena. This illustrates how preliminary explanations in terms of data can be developed into more rigorous analytic theories. Thus, building naturalistic models can be a stepping stone toward more rigorous, normative theories of intelligence as a rational solution to a particular problem setting—much like prior work in cognitive science (e.g. Anderson,, 1991; Oaksford and Chater,, 2009).
@@
These two components serve different purposes. The first helps to demonstrate that the models are real candidate models of the phenomena in question, rather than models that simply could not scale to the real problem or that are overfit to a particular instantiation of the task. This component is necessary because otherwise the problem space is under-constrained. Moreover, it provides a test-bed for experiments that would be impossible in reality, which is useful both scientifically and practically. The second component allows linking these predictions to the explanatory understanding that cognitive scientists typically seek, and that may help us to generalize our insights beyond current paradigms. This perspective aligns with past arguments on how deep learning can contribute to understanding in cognitive neuroscience (Saxe et al.,, 2021; Kanwisher et al.,, 2023; Cao and Yamins, 2024b, ; Doerig et al.,, 2023). However, we emphasize that combining these two components makes explicit links between the actual problem the system solves and the theorized constructs underlying that solution. It also ensures that our theories make testable predictions beyond the narrow settings of a particular task paradigm.
@@ 7 Discussion
In this paper, we have tried to outline a direction of research that we call “naturalistic computational cognitive science,” which aims to build generalizable models of cognition that scale to naturalistic tasks, while still offering routes to explanatory and theoretical understanding. This perspective synthesizes a growing literature on the importance of naturalistic experiments and generalizable models, and grows out of a long-standing focus in cognitive science on explaining a broad range of phenomena. In this section, we outline these connections, and highlight the future directions suggested by our perspective.
@@
The quest for building general models of natural intelligence. A long history of cognitive science research has sought frameworks with the generality to explain a wide scope of cognitive phenomena. Researchers in cognitive architecture (e.g. Ritter et al.,, 2019; Laird,, 2019), connectionist (e.g. McClelland et al.,, 1986, 2003), and Bayesian (e.g. Tenenbaum and Griffiths,, 2001; Griffiths et al.,, 2010; Tenenbaum et al.,, 2011), paradigms have tried to identify underlying principles or mechanisms that explain many phenomena. However, while these modeling frameworks themselves are general, typical instantiations of the frameworks have built specialized models focused on individual tasks. Indeed, one core critiques of many modeling paradigms is their failure to accurately capture patterns of human generalization (e.g. Fodor et al.,, 1975; McClelland and Patterson,, 2002; Lake et al.,, 2017).
@@
In pursuit of generality, some recent studies have built upon foundation models from AI that can perform many tasks. Some works have applied cognitive paradigms to study the behaviors of these models (e.g. Binz and Schulz,, 2023; Lampinen et al.,, 2024; Buschoff et al.,, 2025). Others have finetuned these models using cognitive data, to create generalizable models that can predict behavior on new experimental paradigms in areas like vision (e.g. Muttenthaler et al., 2024b, ; Muttenthaler et al., 2024a, ; Fu et al.,, 2024), or the broader space of cognitive tasks that can be presented in language (Binz and Schulz,, 2024; Binz et al.,, 2024).
@@
We see the arguments that we have laid out in this paper as being broadly compatible with many of these approaches and frameworks. However, we offer a stronger emphasis on the role of theory (cf. Frank and Goodman,, 2025), and the virtuous cycle between expanding the scope and naturalism of our experimental designs, and expanding the generalizability of our models and theories.
@@
Towards experiments that involve naturalistic behavior Our perspective aligns with various works that have likewise highlighted the value of considering complex naturalistic behaviors and the environments in which they occur. The strands of cognitive research focused on embodiment and enactivism have advocated that the environment, and organism-environment interactions, form an important component of cognition (e.g. Clark,, 1998; Varela et al.,, 2017). More generally, naturalistic experiments are frequently advocated in neuroscience (e.g. Mobbs et al.,, 2021; Cisek and Green,, 2024), where even prominent visual processing signals can fundamentally differ in active, naturalistic paradigms (e.g. Amme et al.,, 2024). Similarly, Gao and Ganguli, (2015); Gao et al., (2017) highlight that more complex behavioral paradigms are needed to unlock the full space of neural dynamics, and Nastase et al., (2020) argue that isolating only a few variables in our experiments impairs our ability to understand the full scope of natural behavior. Likewise, Wise et al., (2023) advocate for the need to explore more naturalistic environments and behaviors in reinforcement learning, to grapple with the complexity of behavior in environments closer to the real world. We concur with these works, and emphasize that advances in technology both enable richer experimental paradigms, and the models that can perform them. We differ from many of these works in more strongly emphasizing the role of models that generalize across task paradigms.
@@
Linking between learning models and the brain A large recent literature has explored surprising alignments between the representations learned by task-optimized deep learning models and neural representations (Khaligh-Razavi and Kriegeskorte,, 2014; Yamins et al.,, 2014; Yamins and DiCarlo,, 2016; Schrimpf et al.,, 2021; Sucholutsky et al.,, 2023). These findings have stimulated substantial interest and debate, ranging from critical argument that failures to capture particular phenomena doom these models (e.g. Bowers et al.,, 2023), to suggestions that deep learning models upend the theoretical assumptions of the field (Perconti and Plebe,, 2020; Hasson et al.,, 2020). Other authors have written frameworks that attempt to integrate deep learning models within the existing scientific paradigms (Doerig et al.,, 2023). In this vein, several papers have proposed frameworks for understanding the role of deep learning in cognitive neuroscience (e.g Richards et al.,, 2019; Storrs and Kriegeskorte,, 2019; Cichy and Kaiser,, 2019)—and how these models can contribute to deriving explanatory and theoretical understanding (Cao and Yamins, 2024a, ; Cao and Yamins, 2024b, ; Saxe et al.,, 2021). In this context, Feather et al., (2025) argue for the importance of benchmarks that test alignment of both behavior and representations—and highlight how dataset limitations like lack of stimulus variability can limit our ability to discriminate between models. As above, our perspective aligns with many of these works, but we place greater emphasis on the complementary role of increasing naturalism of experimental paradigms, and learning from the engineering paradigms of AI, rather than simply adopting its models as artifacts.
7.1 Looking forward: challenges & opportunities for naturalistic computational cognitive science
While we have advocated for the importance of naturalism, there are a number of practical and conceptual challenges to achieving these goals. Many of these challenges provide important opportunities for future research to address.
First, as we noted above, identifying the parameters of naturalistic variation that may interact with a particular theoretical construct of interest can be challenging. When we form theories—for object perception, for theory of mind, for how people plan—which “parameters” of the experimental paradigm should we make more naturalistic? In the trolley problem, for example, is it the number of people that must be considered? Is it the way in which people might die? Is it the length of time to decide? While researchers can draw on domain expertise to identify potentially-relevant parameters, there may be unexpected influences from other factors that are not typically manipulated–e.g. the differences between moral thoughts and moral actions when shifting from vignettes to virtual reality (Francis et al.,, 2017). Thus, one important direction for naturalistic research may be to identify more general overarching naturalistic factors—such as perception, embodiment, or action—that tend to influence phenomena across many domains. Ultimately, identifying such overarching factors may help us to identify more general unifying principles of cognition.
Another challenge and opportunity is precisely specifying timescales of learning. Current learning-based theories often conflate evolution, development, and short-term learning into a single process of “optimizing” a model for the task of interest. This is motivated from a rational analysis perspective, where the artifact of interest is the outcome of the optimization process, which serves as a theory for behavior. While this has led to fruitful understanding, it leaves open the question of what exactly is learned during each stage. More complete “learning”-based accounts would specify precisely what an “evolutionary” data distribution would be, and the model behaviors (evolved inductive biases) that are expected from it; a succession of “developmental” data distributions and the model behaviors that would result from them; and finally predict both the behaviors that people begin tasks with and how they will with task exposure. Fortunately, work in this direction has already begun. For example, researchers have begun to explicitly map different timescales of learning to different levels of optimization in a meta-learning process (Wang,, 2021)—and indeed, some works have highlighted how meta-learning can give rise to human-like inductive biases, e.g. for linguistic (McCoy et al.,, 2020; McCoy and Griffiths,, 2025) or compositional (Lake and Baroni,, 2023) generalizations. Thus, a key opportunity for naturalistic computational cognitive science is to develop learning-based account of cognition that begin to specify how learning from different data-distributions within different segments of learning can produce the changing cognitive phenomena we observe throughout ontogeny.
Finally, the approaches we have advocated for introduce practical challenges. The types of modeling and experimental paradigms we discuss here often either push towards or pull from the frontiers of contemporary computer science research. This can require both greater technical expertise and computational resources. While we acknowledge this challenge, we see it as an opportunity for larger-scale collaborative efforts—which are increasingly adopted in computer science research on training large-scale models (e.g. Le Scao et al.,, 2023). By pooling expertise and resources, collaborations make progress at larger scales than individual research groups could. Likewise, as discussed above, collaborations on shared benchmarks (e.g. Schrimpf et al.,, 2020) can allow broader and deeper assessments than any individual researcher could. Furthermore, the growing overlap between cognitive modeling and artificial intelligence brings opportunities for mutually-beneficial collaborations between the two fields. Cognitive scientists have unique expertise in developing evaluations for scientifically inferring what cognitive algorithms may underly behavioral phenomena; on the other hand AI researchers have unique expertise in training models that can perform tasks of interest in naturalistic settings. We hope that this paper will help to encourage deeper collaborations between these fields on their shared questions about the nature and origins of intelligence.
7.2 Conclusions